OpenStudy (anonymous):

Evaluate if x = 2.5 + y = - 1.5

OpenStudy (anonymous):

@whpalmer4 @EclipsedStar @RJAY56 @rabicagrace @ROMAN.ESM @RedNeckOutLaw @Robert136 @RosieF @pooja195 @peachpi @pr16 @Preetha @Preetha @proud_yemeniah_ @prowrestler @greenjetteee @G-Pimp-Squat-Boss @gabbimanges17 @Gambino @ganeshie8 @teddydaddy23

OpenStudy (anonymous):

as given x=2.5+y and also 2.5+y=-1.5 so we can say x=-1.5 from equation two we can find value of y 2.5+y=-1.5 y=-4 so x=-1.5 and y=-4